The value of a "Mint State-63" (unicirclated) 1950 Jefferson nickel minted at Denver is 8/7 the value of a similar condition 1945 nickel minted at Philadelphia. Together, the total value of the two coins is $15.00. What is the value of each coin?
n+8/7n=15

Answer by tutorcecilia(2152)

Using your equation:
n+8/7n=15 [combine like terms]
7/7n+ 8/7n=15
15/7n=15 [solve for the n-term]
(7/15)(15/7)n=(15/1)(7/15)
n=7
So, the Phil coin = n = 7
and the Denver coin = 8/7n=(8/7)(7/1)=8
.
checking:
7+8=15
15=15 [checks out]